OpenStudy (anonymous):

Did I do this Algebra problem correctly? (2a^4b)^3/24b^5 I put a^12/3b^2

OpenStudy (lgbasallote):

yes you did

OpenStudy (lgbasallote):

\[\huge \frac{(2a^4b)^3}{24b^5} \implies \frac{8a^{12} b^3}{24b^5} \implies \frac{a^{12}}{3b^2}\]

OpenStudy (anonymous):

yes u did see here http://www.wolframalpha.com/input/?i=%282a%5E4b%29%5E3%2F24b%5E5

OpenStudy (anonymous):

MOre fun! :) 28m^-4n^6/4mn^-2p^-2 Is it 7n^8/m^5p^2

OpenStudy (lgbasallote):

uhhh doesnt sound right... p^2 should be at the top

OpenStudy (lgbasallote):

\[\huge \frac{28m^{-4} n^6}{4mn^{-2} p^{-2}} \implies \frac{28n^6 n^2 p^2}{4m^4 m}\] see?

OpenStudy (anonymous):

Okay.. 7n^4p^2/m^3 We good? :)

OpenStudy (anonymous):

yes its correct (7 n^4)/(m^3 p^2)

OpenStudy (lgbasallote):

lol i got confused...

OpenStudy (lgbasallote):

m^3 is still wrong

OpenStudy (anonymous):

So... am I write? :P

OpenStudy (anonymous):

by doing this (28×n^6/4×m/n^2)/(m^4 p^2)

OpenStudy (anonymous):

*right

OpenStudy (anonymous):

7n^4p^2/m^3 We good? :)

OpenStudy (lgbasallote):

\[m^4 \times m \ne m^3\]

OpenStudy (lgbasallote):

you're almost right except for that m^3

OpenStudy (lgbasallote):

hint: \[\huge m^4 \times m = m^{4+1}\]

OpenStudy (anonymous):

M^5 :P

OpenStudy (lgbasallote):

right! \[\huge \frac{7n^8 p^2}{m^5}\]

OpenStudy (anonymous):

MOre fun! 20m-4n^5/2mn^-2p^-3 is it 10n^3p^3/m^3

OpenStudy (lgbasallote):

lol you're pushing it arent you =)) anyway if you meant \[\huge \frac{20m^{-4}n^5}{2mn^{-2}p^{-3}}\] then you're wrong

OpenStudy (lgbasallote):

change it into positive exponents first so you wont get confused

OpenStudy (lgbasallote):

your exponents for m and n are wrong

OpenStudy (anonymous):

is n7 and m5?

OpenStudy (anonymous):

And yes, yes I am. :)

OpenStudy (lgbasallote):

yup \[\huge \frac{10n^7 p^3}{m^5}\]

OpenStudy (anonymous):

THANK YOU! :D

OpenStudy (lgbasallote):

welcome :D

OpenStudy (anonymous):

OH my god, this next problem is like, impossible. You ready?

OpenStudy (lgbasallote):

lol lay it on me

OpenStudy (anonymous):

 Volume of sphere = pir^2h